OpenStudy (anonymous):

Which of the following functions shown in the table below could be an exponential function? Check all that apply. x -3 -2 -1 0 1 2 3 f(x) 0.67 1.11 1.82 3 5.95 9.15 14.45 g(x) 0.125 0.25 0.5 1 2 4 8 h(x) 0.25 0.5 0.75 1 1.25 1.5 1.75 k(x) 64 16 4 1 0.25 0.0625 0.0156 A. g(x) B. h(x) C. k(x) D. f(x)

OpenStudy (amoodarya):

k(x) is answer because 64/4=16 16/4=4 4/4=1 1/4=.25 .25/4=.0625 .0625/4=.016525 and g(x) is also because every time is multiplied by 2

OpenStudy (amoodarya):